OpenStudy (anonymous):

The price of a gallon of unleaded gas was $2.81 yesterday. Today, the price rose to $2.87. Find the percentage increase. Round your answer to the nearest tenth of a percent

OpenStudy (anonymous):

please hellllp

OpenStudy (anonymous):

2.1 %

OpenStudy (anonymous):

On Monday, a local hamburger shop sold a combined total of 450 hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Monday?

OpenStudy (anonymous):

Did you understand how I did the last one? Why do you just want the answers? ha...u should try to understand. It would make things a lot easier on yourself. You're going to have to learn how to do problems like this on your own eventually.

OpenStudy (anonymous):

well i'm doing this program for a class i have. i've been working for 3 days (and i mean all day) and it makes you do an assessment test. i'm doing most of them on my own...these are just the ones i can't figure out.

OpenStudy (m):

2.87-2.81 = 2.81x solve for x

OpenStudy (anonymous):

M...that is correct. That's how you figure .021 and then multiplying 100 gives you the percent....do you understand that laurenlynn_07?

OpenStudy (anonymous):

yes I understand

OpenStudy (anonymous):

Okay. Well, for the cheeseburger problem....you know that the number of cheeseburgers plus the number of hamburgers is equal to 450. So, you can express that like this: x=number of cheesburgers and y=number of hamburgers...so, x+y=450...you also no that the number of cheeseburgers in two times as many hamburgers. So., you can write that like this: x=2y So, you have 2 equations that represent the problem: x+y=450 x=2y All you have to do is solve for each variable and and you get your answer. Do you understand that?

OpenStudy (anonymous):

so in other words....(number of cheeseburgers)+2(number of cheeseburgers)=450...combining like terms we get 3(number of cheeseburgers)=450...so the number of cheeseburgers is 150.

OpenStudy (anonymous):

okayyyyy... gotcha. thank you! that really helped me!! :)

OpenStudy (anonymous):

Yep..glad you understand.